# Decentralized Identifier NSFW Not Safe For Work

### v1

@@ 7.1. Feed Generators
We next delve deeper into the content of Feed Generators by analyzing the labels associated by Labelers with the posts. Only 12.6% of Feed Generators have some of their content labeled. We focus on the 0.53% of Feed Generators that have 10% or more of their content labeled and report their most frequent label (Figure 9). Most of these Feed Generators (0.096%) are dedicated to explicit content (e.g. “porn”, “sexual”, “nudity”) or spam. The dominance of such labels is caused by the Bluesky official Labeler that focuses on filtering this kind of content and issues the majority of all the labels. While some independent Labelers issue labels dedicated to describing the content rather filtering it out (e.g. the Final Fantasy Labeler), such use is not yet widespread. Importantly, the flagged content is not removed from the platform. Rather, each user can decide how to react to each specific label.
@@
Figure 10 shows the number of curated posts and likes received by each Feed Generator. We observe that the majority of Feed Generators have a low number of posts and likes. However, a few Feed Generators curate a large number of posts ($>500,000$) or receive a high number of likes ($>16,000$). Interestingly, the number of likes received by Feed Generators is not directly proportional to the number of posts curated by them.
To explore this, we manually investigate the Feed Generators on both extremes. Highly-liked Feed Generators returning no posts are personalized. For instance, “the-algorithm” tailors feeds based on user likes, while “whats-hot” aggregates trending content from a user’s personal network. They do not return any content to “empty” accounts that we use for our crawls. On the other extreme, there are automatic and aggregating Feed Generators. For instance, “siro-feeds-etc” is a Japanese language feed posting hundreds of thousands of an animated celebrity “cyber girl”, while “hebrew-feed” automatically reposts all the content in Hebrew. We also find active and highly-liked Feed Generators that curate their content, potentially manually. This includes content relevant for specific communities such as “blacksky” or “furry-new”.
@@
Gaining Popularity. We also investigate additional factors impacting the popularity of Feed Generators. Figure 11 presents the in-degree (top) and out-degree (bottom) distributions of users based on follow operations. A red shading in both plots indicate the density of accounts that created Feed Generators. The shade of red intensifies as the in-degrees increase and the out-degrees decrease. This indicates that the Feed Generator accounts correspond to users with many followers. Intuitively, popular users are more likely to create Feed Generators and Feed Generators increase the popularity of their creators. We observe a reverse trend for the out-degree though. Users creating Feed Generators follow a small number of other accounts. We calculate the Pearson’s Coefficients for various factors to estimate whether they are predictive of an account attracting followers. We find that the number of Feed Generators created does not correlate with the number of followers on the creator account ($r=$ $0.005$). When calculating the sum of likes on the the Feed Generators created, however, we find correlation at $r=$ $0.533$. This indicates that creating good Feed Generators is a way for users to gain a larger followership.
@@ 7.2. Feed Generator As A service
Skyfeed provides by far the most comprehensive list of features explaining its high market share. For instance, it is the only platform implementing regular expressions. None of the platforms enables feed creators consider information of the feed readers. All the personalized Feed Generators are currently run by their creators. While consisting of only 0.09% of all the Feed Generators, they are among the most popular ones in the network (Figure 10). Adding such features to the Feed Generator As A Service platforms would require a high amount of implementation effort and increases the platform costs. However, their lack makes creating highly-quality, personalized Feed Generators impossible.
@@ 8. Related Work
Several recent works have explored alternative emerging federated social networks. Perhaps closest to our work is an initial investigation highlighting the centralization tendencies within Mastodon (Raman et al., 2019). Since, there have been several related studies of these so-called “fediverse” applications. For instance, (Cava et al., 2021) study the growth of Mastodon, and others have investigated how user behavior differs across server instances (Bin Zia et al., 2024; Cava et al., 2022a, b). One of the main challenges Bluesky attempts to overcome is decentralized content moderation, via its labeling architecture. There have been several studies looking at content moderation labeling in alternative decentralized social network. For instance, Hassan et al. investigate issues with policy implementation in decentralized social networks (Hassan et al., 2021), and Zia et al. (Bin Zia et al., 2022) proposed a federated solution to training moderation models in the fediverse. Bluesky offers an interesting new point on the design space. To the best of our knowledge, we are the first to study this new architecture. Beyond the above federated networks, there have also been studies of various P2P decentralized social networks, perhaps most notably Secure Scuttlebutt (SSB). These offer a P2P event-sharing protocol and an architecture for social applications (Tarr et al., 2019). There are also blockchain-based social networks, such as including memo.cash (Zuo et al., 2023), Steemit (Guidi et al., 2020), and Sapien (Jaatun et al., 2022). Finally, some work focused on assessing centralization of decentralized platforms such as Interplanterary Filesystem (IPFS) (Balduf et al., 2023; Wei et al., 2024).
Our work differs as we focus on an entirely different decentralized architecture: Bluesky. In contrast to prior approaches, Bluesky decomposes social platform functionality in a set of sub-components. These are then opened to competing providers. To the best of our knowledge, we are the first to offer a comprehensive measurement study of Bluesky.
@@ 9. Discussion and conclusion
Ease of use and decentralization. Fully decentralized platforms, despite their benefits, tend to be more difficult to use than centralized ones. Manually handling cryptographic keys or server migrations is too difficult for the majority of users (Whitten and Tygar, 1999). The Bluesky approach tries to strike the right balance. By default, the platform automatically handles key management and DNS domain creation. While giving full control to Bluesky PBS, this procedure looks acceptable for most users and enables easy uptake. At the same time, tech-savvy users can opt for more control by managing their keys and domains themselves. Currently, only a small fraction of users (1.1%) have chosen this option, suggesting that few people wish to exploit this opportunity, or the technical challenges remain too difficult for most to overcome. However, the recent development of alternative services with dedicated support facilitating Bluesky identity management (e.g. NameCheap) might increase this number.
@@
Openness and diversity. Openness translates into diversity when the barrier to entry entry is low. Bluesky has opened a number of components to the community, allowing anybody to implement competitors. Most notably, the content moderation and recommendation sub-components enjoy a diverse and growing ecosystem. There are currently $>$ 40k Feed Generators providing personalized feeds (e.g. the-algorithm), focusing on niche communities (e.g. furry-art) or serving explicit content (e.g. feed-me-porn). While Bluesky PBS hides some of this content from the default view, users can still access them by adjusting their settings. Our results show that the use of content moderation is also growing. We find $62$ Labeler accounts in total, of which $34$ are active. Community-operated Labelers already issue the majority of labels in the network after only two months of their introduction. The freedom to create custom labels and to re-adapt existing ones provides a high level of flexibility. It remains to be seen whether some standardization will be necessary to avoid inconsistencies and misinterpretations as the system grows.
@@
Importantly, Bluesky PBS still controls the Firehose and the AppView, which are arguably the main choke points of the system. This enables deleting user accounts, enforcing rules, and vetting external services. However, it is not clear how a potential migration would work, and the network effect might prevent users from switching to a new service.
@@
Scalability. Maintaining centralized components by Bluesky PBS ensures a good user experience and simplifies the development of federated components. For instance, running an independent Labeler requires only lightweight operations, while the Firehose and AppView handle the heavy lifting of aggregating a global view of the network. However, as the platform grows, these centralized components might become bottlenecks. Based on our measurements, we estimate that the Firehose already outputs $\approx$ 30GB of data per day per subscribed client. In the long run, the platform might need to decentralize these components to scale further.
@@
Legal compliance. Bluesky PBS introduces an inclusive approach to content moderation. While some content is hidden by default, it is still processed and served by the platform. This approach might be problematic if users introduce problematic content such as copyrighted material or child sexual abuse material. Furthermore, the platform uses a git-like structure for storing data. Content can be marked as deleted, but can be still recovered from the repositories. This might be problematic from the perspective of privacy-protecting laws such as GDPR.
@@
Economics. The Bluesky network is currently fueled by the Bluesky PBS, enthusiasts, and early adopters. Our analysis of account and Feed Generator descriptions suggests that multiple users raise money via dedicated services (e.g. patreon.com, ko-fi.com) or point to alternative services hosting the users’ and compensating creators (e.g. youtube.com, twitch.tv). Bluesky PBS does not currently suggest introducing advertisements. However, our analysis of multiple forums and discussions around Bluesky suggests that multiple creators consider introducing ads as posts included in their feed or Feed Generators. In the long run, the ecosystem will require economic incentives to become sustainable and compete with centralized platforms such as Twitter/X that recently started sharing its revenue with content creators (tre, [n. d.]).
@@ Appendix B Ethics
We believe that the benefits of our research significantly outweigh potential harms. This work helps understand the implications of opening and decentralizing social network platforms. We take multiple actions to minimize any potential harm. We exclusively collect publicly available information and follow well established ethical procedures for social data (Townsend and Wallace, 2017). We make no attempt to link activities to other accounts or real-world identifies and the collected datais stored securely within a university silo, and no external access is given. Prior to initiating our scans, we contacted the Bluesky team to agree upon a scanning rate that would not disrupt the normal functioning of their service. Additionally, and to ensure a minimal impact, we implement a solution that exclusively downloads repositories if their content has changed since the last snapshot. We identified issues preventing this originally and upstreamed fixes to the Bluesky open source projects.

### v2

@@ 7.1. Feed Generators
We next delve deeper into the content of Feed Generators by analyzing the labels associated by Labelers with the posts. Only 12.6% of Feed Generators have some of their content labeled. We focus on the 0.53% of Feed Generators that have 10% or more of their content labeled and report their most frequent label (Figure 9). Most of these Feed Generators (0.096%) are dedicated to explicit content (e.g. “porn”, “sexual”, “nudity”) or spam. The dominance of such labels is caused by the Bluesky official Labeler that focuses on filtering this kind of content and issues the majority of all the labels. While some independent Labelers issue labels dedicated to describing the content rather than filtering it out (e.g. the Final Fantasy Labeler), such use is not yet widespread. Importantly, the flagged content is not removed from the platform. Rather, each user can decide how to react to each specific label.
@@
Figure 10 shows the number of curated posts and likes received by each Feed Generator. We observe that the majority of Feed Generators have a low number of posts and likes. However, a few Feed Generators curate a large number of posts ($>400,000$) or receive a high number of likes ($>16,000$). Interestingly, the number of likes received by Feed Generators is not directly proportional to the number of posts curated by them.
To explore this, we manually investigate the Feed Generators on both extremes. Highly-liked Feed Generators returning no posts are personalized. For instance, “the-algorithm” tailors feeds based on user likes, while “whats-hot” aggregates trending content from a user’s personal network. They do not return any content to “empty” accounts that we use for our crawls. On the other extreme, there are automatic and aggregating Feed Generators. For instance, “4dff350a5a3e” is a Japanese language feed tracking hundreds of thousands of posts related to the popular noodle dish “ramen”, while “hebrew-feed” automatically reposts all the content in Hebrew. We also find active and highly-liked Feed Generators that curate their content, potentially manually. This includes content relevant to specific communities such as “blacksky” or “furry-new”.
@@
Gaining Popularity. We also investigate additional factors impacting the popularity of Feed Generators. Figure 11 presents the in-degree (top) and out-degree (bottom) distributions of users based on follow operations. A red shading in both plots indicates the density of accounts that created Feed Generators. The shade of red intensifies as the in-degrees increase and the out-degrees decrease. This indicates that the Feed Generator accounts correspond to users with many followers. Intuitively, popular users are more likely to create Feed Generators and Feed Generators increase the popularity of their creators. We observe a reverse trend for the out-degree though. Users creating Feed Generators follow a small number of other accounts. We calculate the Pearson’s Coefficients for various factors to estimate whether they are predictive of an account attracting followers. We find that the number of Feed Generators created does not correlate with the number of followers on the creator account ($r=$ $0.005$). When calculating the sum of likes on the Feed Generators created, however, we find correlation at $r=$ $0.533$. This indicates that creating good Feed Generators is a way for users to gain a larger followership.
@@ 7.2. Feed Generator As a Service
Skyfeed provides by far the most comprehensive list of features, explaining its high market share. For instance, it is the only platform that supports regular expressions. Additionally, while Skyfeed does support personalized feeds, this feature requires manual setup from the developers and is not automated. As a result, most personalized Feed Generators are currently run by their creators. Although consisting of only 0.09% of all Feed Generators, they are among the most popular ones in the network (Figure 10). Adding such features to the Feed Generator As a Service platforms would require a high amount of implementation effort and increases the platform costs. However, their lack makes creating highly-quality, personalized Feed Generators challenging.
@@
Takeaways. In contrast to Labelers, Feed Generators already play a prominent role in the Bluesky ecosystem. High quality feed generators are widely used by users, while automatic, spamming feed generators are less popular. Running a popular Feed Generator is an efficient way for users to gain new followers providing a potential incentive for users to create them. The already dynamic ecosystem of Feed Generator As a Service platforms and its simplicity of use promises a rapid growth of user-led Feed Generators.
@@ 8. Related Work
Several recent works have explored alternative emerging federated social networks. Perhaps closest to our work is an initial investigation highlighting the centralization tendencies within Mastodon (Raman et al., 2019). Since then, there have been several related studies of these so-called “fediverse” applications. For instance, (Cava et al., 2021) study the growth of Mastodon, and others have investigated how user behavior differs across server instances (Bin Zia et al., 2024; Cava et al., 2022a, b). Additionally, some very recent works deal with Bluesky from a social network perspective, which is not the focus of our work: Quelle and Bovet (Quelle and Bovet, 2024) investigate the social and interaction graphs for Bluesky from a network-scientific perspective. Jeong et al. (Jeong et al., 2024a) collect and investigate a dataset of user interactions with timestamp annotations to find temporal patterns. They also investigate migration patterns from and to Bluesky (Jeong et al., 2024b). One of the main challenges Bluesky attempts to overcome is decentralized content moderation, via its labeling architecture. There have been several studies looking at content moderation labeling in alternative decentralized social networks. For instance, Hassan et al. investigate issues with policy implementation in decentralized social networks (Hassan et al., 2021), and Zia et al. (Bin Zia et al., 2022) proposed a federated solution to training moderation models in the fediverse. Bluesky offers an interesting new point on the design space. To the best of our knowledge, we are the first to study this new architecture. Beyond the above federated networks, there have also been studies of various P2P decentralized social networks, perhaps most notably Secure Scuttlebutt (SSB). These offer a P2P event-sharing protocol and an architecture for social applications (Tarr et al., 2019). There are also blockchain-based social networks, such as including memo.cash (Zuo et al., 2023), Steemit (Guidi et al., 2020), and Sapien (Jaatun et al., 2022). Finally, some work focused on assessing the centralization of decentralized platforms such as Interplanterary Filesystem (IPFS) (Balduf et al., 2023; Wei et al., 2024).
Our work differs as we focus on an entirely different decentralized architecture: Bluesky. In contrast to prior approaches, Bluesky decomposes social platform functionality into a set of sub-components. These are then opened to competing providers. To the best of our knowledge, we are the first to offer a comprehensive measurement study of Bluesky and its novel architecture.
@@ 9. Discussion and conclusion
Ease of use and decentralization. Fully decentralized platforms, despite their benefits, tend to be more difficult to use than centralized ones. Manually handling cryptographic keys or server migrations is too difficult for the majority of users (Whitten and Tygar, 1999). The Bluesky approach tries to strike the right balance. By default, the platform automatically handles key management and DNS domain creation. While giving full control to Bluesky PBC, this procedure looks acceptable for most users and enables easy uptake. At the same time, tech-savvy users can opt for more control by managing their keys and domains themselves. Currently, only a small fraction of users (1.1%) have chosen this option, suggesting that few people wish to exploit this opportunity, or the technical challenges remain too difficult for most to overcome. However, the recent development of alternative services with dedicated support facilitating Bluesky identity management (e.g. NameCheap) might increase this number.
@@
Openness and diversity. Openness translates into diversity when the barrier to entry is low. Bluesky has opened a number of components to the community, allowing anybody to implement competitors. Most notably, the content moderation and recommendation sub-components enjoy a diverse and growing ecosystem. There are currently $>$ 40k Feed Generators providing personalized feeds (e.g. the-algorithm), focusing on niche communities (e.g. furry-art) or serving explicit content (e.g. feed-me-porn). While Bluesky PBC hides some of this content from the default view, users can still access them by adjusting their settings. Our results show that the use of content moderation is also growing. We find $62$ Labeler accounts in total, of which $34$ are active. Community-operated Labelers already issue the majority of labels in the network after only two months of their introduction. The freedom to create custom labels and to re-adapt existing ones provides a high level of flexibility. It remains to be seen whether some standardization will be necessary to avoid inconsistencies and misinterpretations as the system grows.
@@
Importantly, Bluesky PBC still controls the Firehose and the AppView, which are arguably the main choke points of the system. This enables deleting user accounts, enforcing rules, and vetting external services. However, it is not clear how a potential migration would work, and the network effect might prevent users from switching to a new service.
@@
Scalability. Maintaining centralized components by Bluesky PBC ensures a good user experience and simplifies the development of federated components. For instance, running an independent Labeler requires only lightweight operations, while the Firehose and AppView handle the heavy lifting of aggregating a global view of the network. However, as the platform grows, these centralized components might become bottlenecks. Based on our measurements, we estimate that the Firehose already outputs $\approx$ 30GB of data per day per subscribed client. In the long run, the platform might need to decentralize these components to scale further.
@@
Legal compliance. Bluesky PBC introduces an inclusive approach to content moderation. While some content is hidden by default, it is still processed and served by the platform. This approach might be problematic if users introduce problematic content such as copyrighted material or child sexual abuse material. Furthermore, the platform uses a git-like structure for storing data. Content can be marked as deleted, but can be still recovered from the repositories. This might be problematic from the perspective of privacy-protecting laws such as GDPR.
@@
Economics. The Bluesky network is currently fueled by the Bluesky PBC, enthusiasts, and early adopters. Our analysis of account and Feed Generator descriptions suggests that multiple users raise money via dedicated services (e.g. patreon.com, ko-fi.com) or point to alternative services hosting the users’ and compensating creators (e.g. youtube.com, twitch.tv). Bluesky PBC does not currently suggest introducing advertisements. However, our analysis of multiple forums and discussions around Bluesky suggests that multiple creators consider introducing ads as posts included in their feed or Feed Generators. In the long run, the ecosystem will require economic incentives to become sustainable and compete with centralized platforms such as Twitter/X that recently started sharing its revenue with content creators (tre, [n. d.]).
Acknowledgements.
The authors would like to thank the anonymous referees for their valuable comments and helpful suggestions. This work was supported by the Sponsor German Research Foundation (DFG) https://www.dfg.de/ within the Collaborative Research Center (CRC) Grant #SFB 1053: MAKI.
@@ Appendix B Ethics
We believe that the benefits of our research significantly outweigh potential harms. This work helps understand the implications of opening and decentralizing social network platforms. We take multiple actions to minimize any potential harm. We exclusively collect publicly available information and follow well established ethical procedures for social data (Townsend and Wallace, 2017). We make no attempt to link activities to other accounts or real-world identifies and the collected data is stored securely within a university silo, and no external access is given. Prior to initiating our scans, we contacted the Bluesky team to agree upon a scanning rate that would not disrupt the normal functioning of their service. Additionally, and to ensure a minimal impact, we implement a solution that exclusively downloads repositories if their content has changed since the last snapshot. We identified issues preventing this originally and upstreamed fixes to the Bluesky open-source projects.